Solve the linear system by using the method of your choice.
1.8x+3y=3
-2x-2.5y=-5

please tell me which choice you chose, and please explain the steps as best as you can. Thank you!

Answer by checkley77(12844)

1.8x+3y=3 multiply by 2.5
-2x-2.5y=-5 multiply by 3 & add.
4.5x+7.5y=7.5
-6x-7.5=-15
-----------------------
-1.5x=-7.5
x=-7.5/-1.5
x=5 ans.
1.8*5+3y=3
9+3y=3
3y=3-9
3y=-6
y=-6/3
y=-2 ans.
Proof:
-2*5-2.5*-2=-5
-10+5=-5
-5=-5